OpenStudy (anonymous):

9^2n less than equal to 1/6561

OpenStudy (anonymous):

first, draw it.. :)) translate it to a mathematical expression.

OpenStudy (jack1):

9^4 = 6561 so 9^-4 = 1/ 6561 does this help at all?

OpenStudy (callisto):

"9^2n" => 9^2n "less than equal to" => \(\le\) "1/6561" => 1/6561 You know the rest.